### v2

# MolMix: A Simple Yet Effective Baseline for Multimodal Molecular Representation Learning

Abstract
In this work, we propose a simple transformer-based baseline for multimodal molecular representation learning, integrating three distinct modalities: SMILES strings, 2D graph representations, and 3D conformers of molecules. A key aspect of our approach is the aggregation of 3D conformers, allowing the model to account for the fact that molecules can adopt multiple conformations—an important factor for accurate molecular representation. The tokens for each modality are extracted using modality-specific encoders: a transformer for SMILES strings, a message-passing neural network for 2D graphs, and an equivariant neural network for 3D conformers. The flexibility and modularity of this framework enable easy adaptation and replacement of these encoders, making the model highly versatile for different molecular tasks. The extracted tokens are then combined into a unified multimodal sequence, which is processed by a downstream transformer for prediction tasks. To efficiently scale our model for large multimodal datasets, we utilize Flash Attention 2 and bfloat16 precision. Despite its simplicity, our approach achieves state-of-the-art results across multiple datasets, demonstrating its effectiveness as a strong baseline for multimodal molecular representation learning. Our code is publicly available at https://github.com/andreimano/MolMix.
1 Introduction and Related Work
Accurately representing molecular structures is fundamental in computational chemistry and drug discovery [1, 2, 3]. Effective molecular representations enable machine learning models to predict molecular properties, understand chemical behaviors, and accelerate the development of new compounds. Traditional molecular representation methods typically focus on a single modality, such as SMILES strings [4, 5], chemical fingerprints [6], 2D molecular graphs [7], or the 3D geometry of molecules [8, 9]. While effective, these methods overlook important molecular characteristics that can be captured by other modalities [10, 11]. To address this, recent research has introduced multimodal approaches that integrate multiple molecular representations to provide richer representations for machine learning tasks in chemistry. Stärk et al. [12] proposed an information maximization approach to enhance the mutual information between 2D and 3D molecular embeddings. Similarly, Liu et al. [13] used contrastive pre-training to align 2D and 3D representations. Other approaches extract both 2D and 3D features, such as shortest path distances and 3D distance encodings, to build multimodal models [14, 15]. Zhu et al. [16] unified 2D and 3D molecular data in a pre-training framework by predicting either masked 2D structures or 3D conformations. Additionally, language-based models have been integrated with molecular data. Tang et al. [17], Xiao et al. [18], and Srinivas and Runkana [19] leveraged large-scale language models to incorporate textual descriptions of molecules, enhancing molecular property predictions.
While these representations capture certain aspects of molecular structures, they may not fully encompass the variability inherent in molecular conformations. Many molecular properties, such as solubility, toxicity, and binding affinity, are influenced by the range of conformations a molecule can adopt in nature [20]. Utilizing a single geometric representation for a molecule, therefore, restricts the effectiveness of machine learning models. Moreover, identifying which conformers most significantly impact the properties of interest remains difficult. Consequently, creating comprehensive multi-modal representations that integrate multiple 3D conformations continues to be a challenge. To address this, Axelrod and Gómez-Bombarelli [21] propose a scheme where 3D conformer embeddings are first extracted using an equivariant backbone and then aggregated with an attention mechanism, but they do not train on multimodal data. Similarly, Nguyen et al. [22] introduce a conformer aggregation approach that leverages Optimal Transport techniques to obtain a single 3D embedding from multiple conformers, which is subsequently combined with 2D embeddings derived from a GNN. Additionally, Zhu et al. [23] propose MARCEL, a conformer aggregation benchmark alongside models that employ set encoders to pool conformer embeddings with 2D structures and SMILES strings for downstream tasks. These approaches illustrate ongoing efforts to develop more holistic and effective molecular representations by incorporating both 2D and multiple 3D conformational data.
Despite these advances, there remains a need for simple yet effective models that can seamlessly integrate multiple modalities and handle multiple conformers without significant computational overhead. Moreover, recent observations [24, 25] suggest that some model design choices might be unnecessary for strong empirical performance, thereby making the added complexity superfluous and inefficient. To address this challenge, we propose MolMix, a simple yet effective baseline for multimodal molecular representation learning. We employ modality-specific encoders - a transformer for SMILES strings, a GNN for 2D graphs, and equivariant neural networks for 3D conformers - to extract text and node embeddings from each modality. These embeddings are concatenated into a multimodal sequence, separated by special tokens, and fed into a downstream transformer that predicts molecular properties. By leveraging efficient techniques like Flash Attention [26, 27] and bfloat16 precision, MolMix scales to handle large sequences of atom tokens with minimal computational overhead, enabling the direct incorporation of all conformers. Despite its straightforward design, MolMix achieves state-of-the-art results across multiple datasets, demonstrating that simplicity can be highly effective in multimodal molecular representation learning, while the modular design allows us to easily exchange the specific modality encoders.
To summarize, our main contributions are:
  1. 1.
    Simple multimodal molecular framework: We introduce MolMix, which seamlessly combines SMILES strings, 2D molecular graphs, and multiple 3D conformers into a unified sequence for molecular representation learning.
  2. 2.
    Conformer aggregation: By incorporating node embeddings from 3D conformers, MolMix effectively captures conformational variability.
  3. 3.
    Scalability: We utilize Flash Attention and bfloat16 (bf16) precision to scale our model, enabling the processing of large multimodal datasets with minimal computational overhead.
  4. 4.
    State-of-the-Art performance: MolMix achieves superior results on multiple benchmark datasets, establishing a strong baseline for future research in multimodal molecular representation learning.
  5. 5.
    Transfer learning capabilities We show that MolMix could potentially be used for pre-training on large molecular datasets.
We make our code publicly available at https://github.com/andreimano/MolMix.
2 MolMix: A Multimodal Molecular Transformer
1D Encoder
We represent molecules using SMILES strings, which encode chemical structures as sequences of characters. Let $S=[s_{1},s_{2},\dots,s_{n}]$ denote the input SMILES string, where each $s_{i}$ is a character. Each $s_{i}$ is mapped to an embedding vector $\mathbf{e}_{i}=\text{Embedding}(s_{i})$. To account for sequence order, positional encodings are added: $\mathbf{z}_{i}=\mathbf{e}_{i}+\text{PE}(i)$. A transformer encoder [28] then processes these vectors to obtain the hidden representations
$\mathbf{h}_{i}^{\text{1D}}=\text{Transformer}(\mathbf{z}_{i}),$ (1)
for all $i\in\{1,\dots,n\}$. Each hidden representation $\mathbf{h}_{i}^{\text{1D}}$ corresponds to the respective input character $s_{i}$, effectively capturing the contextual information about the molecule, for each character.
2D Encoder
We represent molecules as graphs $G=(V,E)$, where $V$ is the set of atoms and $E$ is the set of covalent bonds. Each atom $v\in V$ and bond $e_{uv}\in E$ are associated with initial feature vectors $\mathbf{x}_{v}$ and $\mathbf{e}_{uv}$, respectively. We use a message-passing framework with GINE [29, 30] as the backbone to capture the molecular graph’s structural information. At each message-passing step $j$, the hidden representation of atom $v$ is updated as
$\mathbf{h}_{v,j}^{\text{2D}}=\text{GINE}\left(\mathbf{h}_{v,j-1}^{\text{2D}},\{\mathbf{h}_{u,j-1}^{\text{2D}}\mid u\in\mathcal{N}(v)\},\{\mathbf{e}_{uv}\}\right),$ (2)
where $\mathcal{N}(v)$ denotes the neighbors of atom $v$. This iterative process aggregates information from neighboring atoms and bonds, enabling the model to learn graph representations. The final hidden embeddings $\mathbf{h}_{v,j}^{\text{2D}}$ encode both local and global structural features of the molecule.
3D Encoder
To leverage the three-dimensional structural information of molecules, we utilize 3D conformations represented by the spatial coordinates of each atom. Let $V$ denote the set of atoms. Each atom $v\in V$ is associated with a 3D coordinate $\mathbf{r}_{v}\in\mathbb{R}^{3}$. To extract meaningful atom embeddings that respect the geometric properties of the molecule, we employ an neural network with 3D inductive biases, such as SchNet [8] or GemNet [9], as the backbone model. These networks process the 3D coordinates $\{\mathbf{r}_{v}\}_{v\in V}$ along with the initial atom features $\{x_{v}\}_{v\in V}$ and apply a cutoff function to consider interactions within a specified distance range, generating hidden embeddings
$\mathbf{h}_{v}^{\text{3D}}=\text{3DNetwork}(\mathbf{r}_{v},\mathbf{x}_{v}),$ (3)
for all $v\in V$. These atom embeddings $\mathbf{h}_{v}^{\text{3D}}$ capture both the local geometry and the global spatial arrangement of the molecule.
Downstream Multimodal Transformer
To integrate different molecular representations, we design a multimodal transformer that combines three distinct modalities. The SMILES encoder outputs token embeddings $\mathbf{h}_{i}^{\text{1D}}$, where $\mathbf{h}_{i}^{\text{1D}}$ corresponds to the $i^{\text{th}}$ character in the string. From the 2D MPNN encoder, we extract atom embeddings $\mathbf{h}_{v,j}^{\text{2D}}$ for atom $v$ at layer $j$. By using embeddings from all layers, the model captures both local and distant atom interactions, mitigating the oversmoothing effect common in deep GNNs. The 3D encoder provides atom embeddings $\mathbf{h}_{v,c}^{\text{3D}}$ for atom $v$ and conformer $c$, encapsulating spatial geometry. We use multiple conformers by simply adding all the atom embeddings to the multimodal sequence. Modality-specific learnable encodings are added to the embeddings from each modality. These modality-enhanced embeddings are concatenated into a unified sequence, with special tokens included: a classification token $\mathbf{h}_{\text{CLS}}$ is added at the start, and separation tokens $\mathbf{h}_{\text{SEP}}$ are placed between modalities. The resulting input sequence is structured as
$\mathbf{H}=\left[\mathbf{h}_{\text{CLS}},\{\mathbf{h}_{i}^{\text{1D}}\}_{i},\mathbf{h}_{\text{SEP}},\left\{\mathbf{h}_{v,j}^{\text{2D}}\right\}_{v,j},\mathbf{h}_{\text{SEP}},\left\{\mathbf{h}_{v,c}^{\text{3D}}\right\}_{v,c},\mathbf{h}_{\text{SEP}}\right].$
This sequence is then processed by the downstream transformer, which utilizes the self-attention mechanism to integrate and contextualize information across all modalities. After the transformer layers, the embedding corresponding to the classification token $\mathbf{h}^{out}_{CLS}$ is extracted and sent to a readout MLP to perform downstream tasks such as property prediction and molecular classification:
[…]
$\hat{y}$ $=\text{MLP}(\mathbf{h}^{\text{out}}_{\text{CLS}})$ (5)
We reduce memory overhead in our multimodal transformer with bfloat16 precision and Flash Attention 2 [27]. See appendix C for details and a memory comparison with classical attention.
Since we use the same positional encoding for each $\mathbf{h}_{v,i}^{\text{2D}}$ and $\mathbf{h}_{v,c}^{\text{3D}}$, we maintain the permutation equivariance property of the 2D and 3D encoders. Another desirable property is for the model to preserve any invariance of the 3D encoder. Indeed, MolMix preserves these useful inductive biases:
Theorem 1.
Let $S$ be the SMILES string, $G$ be the 2D graph, and $\{c_{1},\dots,c_{k}\}$ be a set of $k$ 3D conformers for a molecule. Let $\hat{y}=f_{\theta}(S,G,\{c_{1},\dots,c_{k}\})$ be the output prediction obtained as described in eq. 1 - (5). Let our 3D encoder be invariant to the actions of some group $\mathcal{G}$. Then $f_{\theta}$ is also invariant to any $T_{1},\dots,T_{k}\in\mathcal{G}$, i.e. $f_{\theta}(S,G,\{T_{1}c_{1},\dots,T_{k}c_{k}\})=f_{\theta}(S,G,\{c_{1},\dots,c_{k}\})$.
Table 1: Comparison between MolMix and other baselines on the Drugs-75K and Kraken datasets from MARCEL [23]. 1D, 2D and 3D represents training on the SMILES strings and molecule fingerprints, 2D molecular representations and 3D conformers. Multimodal represents training on all three modalities. The metric used is the Mean Absolute Error (MAE, $\downarrow$). Bold indicates the best-performing model, while underline denotes the second-best. MolMix obtains the best results on 5 out of 7 properties, with second-best results obtained on two properties (Drugs-75K/$\chi$ and Kraken/BurL).
Model Drugs-75K Kraken
IP $\downarrow$ EA $\downarrow$ $\chi$downarrow B5 $\downarrow$ L $\downarrow$ BurB5 $\downarrow$ BurL $\downarrow$
1D RF 0.498±0.003 0.474±0.002 0.273±0.003 0.476±0.004 0.430±0.009 0.275±0.018 0.152±0.014
LSTM 0.478±0.002 0.464±0.000 0.250±0.005 0.487±0.028 0.514±0.041 0.281±0.004 0.192±0.002
Transformer 0.661±0.002 0.585±0.003 0.407±0.001 0.961±0.081 0.839±0.043 0.493±0.037 0.278±0.021
2D GIN 0.435±0.003 0.417±0.003 0.226±0.002 0.313±0.026 0.400±0.034 0.172±0.003 0.120±0.004
GIN+VN 0.436±0.006 0.417±0.008 0.227±0.000 0.357±0.003 0.434±0.042 0.242±0.003 0.174±0.011
ChemProp 0.460±0.003 0.442±0.005 0.244±0.001 0.485±0.007 0.545±0.045 0.300±0.009 0.195±0.014
GraphGPS 0.435±0.005 0.409±0.006 0.221±0.005 0.345±0.032 0.436±0.013 0.207±0.012 0.150±0.014
3D SchNet 0.439±0.006 0.421±0.002 0.224±0.009 0.329±0.007 0.546±0.034 0.230±0.011 0.186±0.010
DimeNet++ 0.444±0.009 0.423±0.007 0.244±0.008 0.351±0.011 0.417±0.040 0.210±0.016 0.153±0.007
GemNet 0.407±0.001 0.392±0.002 0.197±0.004 0.279±0.013 0.375±0.009 0.178±0.010 0.164±0.006
PaiNN 0.451±0.004 0.450±0.005 0.232±0.004 0.344±0.039 0.447±0.032 0.240±0.018 0.167±0.009
ClofNet 0.439±0.008 0.425±0.007 0.238±0.002 0.487±0.009 0.642±0.036 0.288±0.017 0.253±0.005
LEFTNet 0.417±0.001 0.396±0.001 0.208±0.005 0.307±0.001 0.449±0.026 0.218±0.001 0.149±0.010
Multimodal SchNet 0.454±0.007 0.438±0.013 0.237±0.010 0.270±0.019 0.432±0.046 0.202±0.018 0.144±0.004
DimeNet++ 0.413±0.008 0.394±0.003 0.227±0.005 0.263±0.012 0.347±0.009 0.178±0.011 0.119±0.011
GemNet 0.419±0.002 0.400±0.001 0.217±0.004 0.231±0.003 0.339±0.027 0.159±0.007 0.095±0.001
PaiNN 0.447±0.007 0.427±0.003 0.229±0.007 0.223±0.022 0.362±0.019 0.169±0.011 0.132±0.009
ClofNet 0.428±0.006 0.403±0.002 0.220±0.007 0.323±0.002 0.449±0.005 0.218±0.019 0.155±0.004
LEFTNet 0.417±0.004 0.395±0.000 0.207±0.002 0.264±0.013 0.364±0.035 0.202±0.003 0.139±0.001
MolMix 0.405±0.002 0.379±0.004 0.206±0.002 0.191±0.017 0.305±0.020 0.146±0.002 0.121±0.005
3 Experimental Setup and Results
Table 2: Comparison between MolMix and other approaches as reported in [22]. The metric used is Root Mean Squared Error (RMSE $\downarrow$). MolMix obtains the overall best scores, significantly improving upon the recently proposed multimodal CONAN-FGW model.
Model Lipo $\downarrow$ ESOL $\downarrow$ FreeSolv $\downarrow$ BACE $\downarrow$
2D-GAT 1.178±0.454 1.513±0.130 2.926±1.160 1.358±0.574
D-MPNN 0.731±0.148 0.961±0.212 2.053±0.261 0.850±0.145
MolFormer 0.701±0.110 0.875±0.249 2.342±0.212 1.045±0.145
SchNet-scalar 0.839±0.179 0.820±0.164 1.268±0.397 0.850±0.316
SchNet-emb 0.767±0.179 0.797±0.239 1.260±0.369 0.832±0.167
ChemProp3D 0.695±0.230 0.825±0.152 1.419±0.427 0.903±0.412
CONAN 0.746±0.114 0.756±0.138 1.223±0.397 0.797±0.226
CONAN-FGW 0.650±0.126 0.727±0.148 1.033±0.288 0.741±0.126
MolMix 0.614±0.022 0.639±0.017 0.976±0.044 0.387±0.041
In this section, we evaluate how MolMix improves predictive performance on real-world datasets by addressing the following questions: Q1) How does MolMix’s performance compare to other sophisticated models?; Q2) Does incorporating multiple modalities enhance downstream performance?; Q3) Are pre-trained weights beneficial for transfer learning?
To address Question 1, we train MolMix on four MoleculeNet datasets [31]Lipo, ESOL, FreeSolv, and BACE—covering various molecular properties, including physical chemistry and biophysics. Conformers are generated using the RDKit chemoinformatics package [32]. We follow the same train/validation/test splits as [22]. We also train models on the newly introduced MARCEL benchmark [23], specifically the Drugs-75k and Kraken datasets. Drugs-75K, a subset of the GEOM-Drugs dataset [33], contains 75,099 molecules with conformers generated by Auto3D [34], and labels for ionization potentials (IP), electron affinity (EA), and electronegativity ($\chi$). Kraken [35] includes 1,552 monodentate organophosphorus ligands, with conformers generated via DFT, and labels for four 3D ligand descriptors: Sterimol B5 (B5), Sterimol L (L), buried Sterimol B5 (BurB5), and buried Sterimol L (BurL). We follow the same splits as in [23].
Figure 1: Modality ablation study on the Kraken dataset (MAE $\downarrow$). We keep the downstream Transformer fixed and train using a single modality or a combination of modalities. Using all three modalities obtains the best results on three out of the four properties, with the second-best results generally being obtained by a configuration that contains 3D conformers. Notably, for the buried Sterimol L property, the best results are obtained by a 3D encoder + Transformer model, indicating that the property could mainly depend on the 3D structure.
Modality B5 $\downarrow$ L $\downarrow$ BurB5 $\downarrow$ BurL $\downarrow$
1D 0.499±0.033 0.497±0.025 0.291±0.020 0.187±0.005
2D 0.258±0.017 0.347±0.013 0.176±0.010 0.141±0.004
3D 0.213±0.008 0.337±0.009 0.164±0.004 0.116±0.003
1D+2D 0.297±0.015 0.390±0.016 0.180±0.008 0.153±0.006
1D+3D 0.209±0.002 0.337±0.010 0.156±0.013 0.127±0.004
2D+3D 0.202±0.009 0.356±0.026 0.151±0.004 0.122±0.004
1D+2D+3D 0.191±0.017 0.305±0.020 0.146±0.002 0.121±0.005
Figure 2: Transfer learning experiment. We select the best checkpoint of a model trained to predict the electronegativity ($\chi$) on the Drugs-75K dataset. We then freeze the model and only train the last linear readout layer on the Kraken dataset. We compare with a randomly initialized model. For all descriptors, using the pre-trained weights improve predictive performance. Note that pretraining improves both mean performance and standard deviations.
Modality B5 $\downarrow$ L $\downarrow$ BurB5 $\downarrow$ BurL $\downarrow$
Random init. 0.567±0.010 0.543±0.020 0.334±0.004 0.216±0.003
Pretrain 0.521±0.003 0.509±0.004 0.316±0.001 0.195±0.003
For all experiments, we report the mean and standard deviation over five different runs. We use the same hyperparameters for the modality encoders—we encode the SMILES strings using a transformer with two layers and four attention heads, the 2D graph is processed by a GINE [29] network containing 6 layers, and the 3D conformations are processed either using a SchNet [8] or a GemNet [9] model. All of the encoders have a hidden dimension of 128. The modality embeddings are then projected by a linear layer to a 512-dimensional space before we jointly processing them with a downstream Transformer network with 8 heads and 6 layers for the Lipo, ESOL, FreeSolv, BACE and Kraken, and 12 heads with 8 layers on Drugs-75K. We use the Schedule-Free AdamW optimizer [36, 37, 38].
On the MoleculeNet datasets, MolMix obtains the overall best results, significantly improving upon the results of [22] on some datasets such as BACE, as can be seen in table 2. This highlights that our simple approach can potentially learn better multimodal representations with conformer aggregation than previously proposed methods that contain more sophisticated aggregation techniques.
On the MARCEL datasets, compared to the approaches in [23], we achieve the best results on five out of seven properties and second-best on the remaining two, as can be seen in table 1. This suggests that MolMix also consistently performs well when using physically-grounded conformer generation methods like DFT and Auto3D.
To address Question 2, fig. 2 shows that for three of the four Kraken descriptors, training on all three modalities yields the best results, while for one property, training on the 3D modality alone performs slightly better. The results indicate that 3D tokens contribute the most to downstream performance, followed by 2D, with SMILES (1D) having the least impact. Notably, SMILES strings significantly improve performance when predicting the Sterimol L descriptor.
To answer Question 3, we select the best checkpoint from a model trained to predict electronegativity ($\chi$) on the Drugs-75K dataset. We then freeze its weights and train only the final linear layer to predict descriptors on the Kraken dataset, comparing it to a randomly-initialized model. As shown in fig. 2, pre-training improves predictive performance in all cases, suggesting that with sufficient data, MolMix could serve as a foundation model for molecular tasks.
4 Conclusions and Further Work
We propose MolMix, a simple yet effective multimodal molecular transformer supporting conformer aggregation. MolMix preserves inductive biases of modality encoders and achieves state-of-the-art results across multiple datasets. We hint towards MolMix being able to support transfer learning, suggesting that it could be used as a molecular foundation model. Finally, we use Flash Attention and bf16 precision to handle longer sequences and multiple modalities efficiently.
We leave three open questions. First, large self-supervised VLMs excel in 0-shot prediction and fine-tuning [39, 40, 41, 42]. Exploring self-supervised pre-training for MolMix using signals like masked language modeling [43] and noise-contrastive estimation [44] could be valuable. Second, multiple conformers without pooling may be suboptimal; token merging [45] could improve memory and runtime. Lastly, adding modalities like molecular fingerprints may enhance performance.
Acknowledgments
[…]
Appendix A Proof for Theorem 1
Theorem 1.
Let $S$ be the SMILES string, $G$ be the 2D graph, and $\{c_{1},\dots,c_{k}\}$ be a set of $k$ 3D conformers for a molecule. Let $\hat{y}=f_{\theta}(S,G,\{c_{1},\dots,c_{k}\})$ be the output prediction obtained as described in eq. 1 - (5). Let our 3D encoder be invariant to the actions of some group $\mathcal{G}$. Then $f_{\theta}$ is also invariant to any $T_{1},\dots,T_{k}\in\mathcal{G}$, i.e. $f_{\theta}(S,G,\{T_{1}c_{1},\dots,T_{k}c_{k}\})=f_{\theta}(S,G,\{c_{1},\dots,c_{k}\})$.
Proof.
Let $g_{\theta}$ be our 3D encoder network, as described in eq. 3. Let $V$ be the set of atoms and $\{x_{v}\}_{v\in V}$, $\{r_{v}\}_{v\in V}$ the atom features and their 3D coordinates, such that a conformer can be described as the tuple $c=(\{x_{v}\}_{v\in V},\{r_{v}\}_{v\in V})$. We assume that $g_{\theta}$ is invariant to any action $T\in\mathcal{G}$, therefore we have that, for any conformer $c$, $g_{\theta}(Tc)=g_{\theta}(c)=\mathbf{h}^{\text{3D}}$.
Let $h_{\theta}$ be the downstream Transformer together with the readout layer, as described in eq. 4 - (5). Since we add the same learnable modality encoding to each $h^{\text{3D}}_{v,k}$, we also have that for any permutation $\pi\in Sym(K)$, we have
$h_{\theta}(\{g_{\theta}(T_{1}c_{1}),\dots,g_{\theta}(T_{k}c_{k})\})$ $=h_{\theta}(\{g_{\theta}(c_{1}),\dots,g_{\theta}(c_{k})\})$
$=h_{\theta}(\{\mathbf{h}_{v,1}^{\text{3D}},\dots,\mathbf{h}_{v,k}^{\text{3D}}\}_{v\in V})$
$=h_{\theta}(\{\mathbf{h}_{v,\pi(1)}^{\text{3D}},\dots,\mathbf{h}_{v,\pi(k)}^{\text{3D}}\}_{v\in V})$
$=\tilde{y},$
therefore, if when we include the 2D graph $G$ and the SMILES string $S$, we obtain $f_{\theta}(S,G,\{T_{1}c_{1},\dots,T_{k}c_{k}\})=f_{\theta}(S,G,\{c_{1},\dots,c_{k}\})=\hat{y}$. ∎
Appendix B Qualitative attention example
Figure 3: Red lines mark the boundaries between modalities on the key axis (with keys for each token represented by columns), while yellow lines mark the boundaries on the query axis (with queries represented by rows). The modalities are ordered as 3D, SMILES, and 2D. The attention scores are taken from the first layer of the model and clipped to the $[-10,10]$ range.
We present the attention scores for each head in a MolMix model trained on the Drugs-75k dataset, using a randomly sampled molecule from the dataset for visualization. As shown in Figure 3, distinct patterns emerge across the attention heads. While it remains challenging to assign a definitive interpretation to each individual head, certain sparse or dense patterns are evident in each cross-modality section. This suggests that the model is learning to extract meaningful and potentially useful features from all modalities.
Appendix C Attention implementation details
We employ Flash Attention 2 [27] for the self-attention mechanism in our models. Flash Attention 2 is a hardware-optimized implementation that significantly reduces both memory usage and runtime compared to the standard attention algorithm. It achieves these gains by leveraging GPU programming techniques, such as kernel fusion and tiling. Additionally, we utilize the varlen implementation, which prevents unnecessary memory and compute consumption on padding tokens.
Table 3 presents the memory savings achieved by using Flash Attention 2 during training.
Table 3: Comparison between fp32 standard attention and bf16 Flash Attention 2 memory usage across different models and batch sizes.
Dataset Kraken
Batch Size 16 32 64
Standard Attention (SchNet)  67 GB OOM OOM
Standard Attention (GemNet)  80 GB OOM OOM
Flash Attention 2 (SchNet)  5.1 GB  11.6 GB  22.3 GB
Flash Attention 2 (GemNet)  22 GB  40 GB  73 GB